# Urban Region Representation Learning: A Flexible Approach

### v1

@@ Appendix I Details of Baseline Models
Following HAFusion, we also evaluate the applicability of our model to areas of different urban environments using regions in Staten Island, which is the largest borough of New York City by land area yet the least densely populated. This contrasts the area of the NYC dataset used above that covers Manhattan, which is the smallest borough by land area but the most densely populated.
@@
We report in Fig. 10 the $R^{2}$ results for the regions in these two boroughs, focusing on the check-in count prediction task for conciseness. All models report lower accuracy (i.e., smaller $R^{2}$) on Staten Island than on Manhattan. This is expected, as Staten Island is less densely populated with smaller variations in the urban features to help the models learn distinctive embeddings for different regions. The mobility data-based models (MVURE, MGFN, HREP, ReCP, and HAFusion) suffer the most, as the mobility data become more sparse. For example, Manhattan has about ten million records in a single month, whereas Staten Island only has hundreds.
FlexiReg again outperforms all competitors, now with an even larger performance gap. The model does not rely on mobility data, while its adaptive aggregation module enables it to adaptively fuse different input features of grid cells to accommodate areas of different urban characteristics.
Figure 10. Model applicability to suburban areas.
M.3. Impact of Grid Cell Design
Figure 11. Impact of grid cell design.
@@
As Fig. 11 shows, FlexiReg produces the best (i.e., largest) $R^{2}$ scores across all four downstream tasks, confirming the effectiveness of our default grid cell design.
@@
FlexiReg-LargeHex has the worst accuracy – it has only 89 cells while FlexiReg has 438. The larger cells can mix the distinctive features in a cell, missing local urban characteristics. When the cells become larger than the target regions for the downstream tasks, capturing the distinctive urban features within a region becomes even more challenging.
In contrast, FlexiReg-SmallHex has 3,201 cells. Now each cell is too small, such that urban features become sparse in each cell, which makes it difficult to learn meaningful embeddings. These smaller cells also intrigue additional computational costs, requiring more processing times (57, 208, and 1294 seconds for FlexiReg-SmallHex, FlexiReg, and FlexiReg-LargeHex, respectively). These findings ground our design choice of the hexagonal grid.
M.4. Impact of Model Parameter Values
@@
Impact of the region embedding dimensionality $d$. We start by varying $d$ from 36 to 288. We test all models except for UrbanCLIP and CityFM whose $d$ cannot be easily varied as discussed in Appendix J. The learned region embeddings are used for the four downstream prediction tasks as before.
@@
Impact of the number of street view images per region used in SV-RAlign, #$\#SV$. The number of street view images per region used in SV-RAlign reflects the richness of ground-level visual patterns captured. We vary #$\#SV$ from 16 to 256 to evaluate its impact on the performance of FlexiReg. As Table 11 shows, FlexiReg achieves the best performance with different values of #$\#SV$ on different downstream tasks. The variation in $R^{2}$ for different #$\#SV$ values on the same task is marginal, against verifying the robustness of our model. A larger #$\#SV$ value, i.e., using more street view images, allows the model to capture more comprehensive visual patterns. It takes more time to process the images, and it could introduce noisy and conflicting patterns. We have set #$\#SV$ as 64 by default to balance between learning effectiveness and efficiency.
@@
Impact of the pre-trained CNN model $CNN_{PT}$ used in the CNN branch. The pre-trained CNN model impacts the capability of extracting features from satellite images. We explore four different pre-trained CNN models (AlexNet (Krizhevsky et al., 2012), EfficientNet (Tan and Le, 2019), ResNet-18 (He et al., 2016), and ResNet-34 (He et al., 2016)) for our model. For a clearer comparison over the impact of the use of these CNN models, we remove the prompt enhancer module from FlexiReg and focus solely on the multi-modal grid cell embedding learning module.

### v3

@@ 4.7. Applicability to Suburban Areas (Q5)
We also evaluate the applicability of our model to areas of different urban environments using regions in Staten Island, which is the largest borough of New York City by land area yet the least densely populated. This contrasts the area of the NYC dataset used above that covers Manhattan, which is the smallest borough by land area but the most densely populated.
Figure 10. Model applicability to suburban areas (NYC).
@@
We report in Fig. 10 the $R^{2}$ results for the regions in these two boroughs, focusing on the check-in count prediction task for conciseness. All models report lower accuracy (i.e., smaller $R^{2}$) on Staten Island than on Manhattan. This is expected, as Staten Island is less densely populated with smaller variations in the urban features to help the models learn distinctive embeddings for different regions. The mobility data-based models suffer the most, as the mobility data become more sparse. FlexiReg again outperforms all competitors, now with an even larger performance gap. The model does not rely on mobility data, while its adaptive aggregation module enables it to adaptively fuse different input features of grid cells to accommodate areas of different urban characteristics.
4.8. Impact of Grid Cell Design (Q6)
@@
As Table 10 shows, FlexiReg produces the best (i.e., largest) $R^{2}$ scores across all four downstream tasks, confirming the effectiveness of our default grid cell design.
Table 10. Impact of grid cell design ($R^{2}\uparrow$ on NYC)
Crime Check-in Service Call Population
FlexiReg-Rect 0.741 0.836 0.557 0.699
FlexiReg-LargeHex 0.587 0.784 0.425 0.233
FlexiReg-SmallHex 0.709 0.803 0.586 0.710
FlexiReg 0.798 0.876 0.601 0.715
@@
FlexiReg-LargeHex has the worst accuracy – it has only 89 cells while FlexiReg has 438. The larger cells can mix the distinctive features in a cell, missing local urban characteristics. When the cells become larger than the target regions, capturing the distinctive urban features within a region becomes even more challenging. In contrast, FlexiReg-SmallHex has 3,201 cells. Now each cell is too small, and urban features become sparse, which makes it difficult to learn meaningful embeddings. These smaller cells also incur more processing times (57, 208, and 1294 seconds for FlexiReg-SmallHex, FlexiReg, and FlexiReg-LargeHex, respectively). These findings ground our choice of using a hexagonal grid.
4.9. Impact of Prompt Templates (Q6)
As Fig. 4 shows, we use a simple prompt template that contains only a task instruction (i.e., the first sentence) which will be populated with information specific to each related cell (e.g., its address).
Table 11. Impact of prompt templates ($R^{2}\uparrow$ on NYC)
Crime Check-in Service Call Population
FlexiReg-ReM 0.786 0.872 0.598 0.709
FlexiReg-RePh 0.793 0.869 0.600 0.710
FlexiReg-RePos 0.788 0.872 0.603 0.712
FlexiReg 0.789 0.876 0.601 0.715
We study the impact of prompt templates with three variants: (1) FlexiReg-ReM removes the task instruction; (2) FlexiReg-RePh rephrases the task instruction with ChatGPT; (3) FlexiReg-RePos moves the task instruction to the end.
We repeat the embedding learning and downstream task prediction tasks as above. As Table 11 shows, the model variants achieve similar accuracy to FlexiReg, indicating that our model’s effectiveness does not depend on a specific prompt template and is robust to prompt engineering. This robustness stems from our use of last-layer token embeddings from the LLM as textual embeddings, avoiding reliance on the generated text. Additionally, our proposed T-RAlign module effectively extracts task-relevant information from these embeddings.
4.10. Impact of Model Parameter Values (Q6)
@@
4.10.1. Impact of the region embedding dimensionality $d$.
We start by varying $d$ from 36 to 288. We test all models except for UrbanCLIP and CityFM whose $d$ cannot be easily varied as discussed in Section 4.1. The learned region embeddings are used for the four downstream prediction tasks as before.
@@
4.10.2. Impact of the text-region embedding dimensionality $d_{text}$.
@@
4.10.3. Impact of the number of street view images per region used in SV-RAlign, #$\#SV$.
The number of street view images per region used in SV-RAlign reflects the richness of ground-level visual patterns captured. We vary #$\#SV$ from 16 to 256 to evaluate its impact on the performance of FlexiReg. As Table 13 shows, FlexiReg achieves the best performance with different values of #$\#SV$ on different downstream tasks. The variation in $R^{2}$ for different #$\#SV$ values on the same task is marginal, against verifying the robustness of our model. A larger #$\#SV$ value, i.e., using more street view images, allows the model to capture more comprehensive visual patterns. It takes more time to process the images, and it could introduce noisy and conflicting patterns. We have set #$\#SV$ as 64 by default to balance between learning effectiveness and efficiency.
4.10.4. Impact of the pre-trained CNN model $CNN_{PT}$ used in the CNN branch.
The pre-trained CNN model impacts the capability of extracting features from satellite images. We explore four different pre-trained CNN models (AlexNet (Krizhevsky et al., 2012), EfficientNet (Tan and Le, 2019), ResNet-18 (He et al., 2016), and ResNet-34 (He et al., 2016)) for our model. For a clearer comparison over the impact of the use of these CNN models, we remove the prompt enhancer module from FlexiReg and focus solely on the multi-modal grid cell embedding learning module.
@@
5. Related Work
These existing studies primarily focus on the general region embedding learning stage, while paying limited attention to the downstream task learning phase. Prompting learning provides a promising solution for downstream task-based learning, which can guide the general region embeddings to adapt to specific tasks. This approach has already achieved significant success in fields such as natural language processing (Schick and Schütze, 2021; Ben-David et al., 2022; Han et al., 2022) and computer vision (Zhou et al., 2022a, b; Khattak et al., 2023). Recently, prompt learning has also been introduced into the field of urban computing (Zhang et al., 2023c; Yuan et al., 2024b). However, the urban tasks targeted in these works differ in focus from ours.
In the context of region embedding learning, only one prior attempt exists: HREP [62], which generates prompt embeddings randomly, without conditioning on any input features. While the prompt embeddings may be correlated with the downstream tasks, they fail to capture the correlation between the region features and downstream tasks.
Discussion. Existing high-performing models rely heavily on human mobility data, while accessible features and their relevance to downstream tasks have been underexplored. We propose a novel, multimodal grid cell embedding learning module and an environment context-based contrastive learning approach to capture distinctive environmental characteristics and spatial correlations between different types of features. Besides, we introduce a prompt enhancer module to extract and integrate task-specific information into generic region embeddings to tailor for diverse tasks.
6. Conclusion